# RadarLLM: Adapting Pretrained Large Language Models for Marine Radar Target Detection with Preference-aware Loss

### v1

@@ III-E2 Theoretical Insight of the Effectiveness of Feature Token Importance
III-F Stage 4: Retrain Classification Head Based on Autoencoder Network
As illustrated in Fig. 1, the output of the LLM-based model has a shape of $[B,K,2]$, representing predictions across multiple tokens. This token-level output cannot be directly used for detection due to the absence of a unified decision representation. To address this, we propose retraining a classifier based on the learned representation $F^{\mathrm{LN}}$. Specifically, we employ an autoencoder-based classification head designed to extract the most informative features from the high-dimensional input, thereby enabling effective and robust decision-making. The proposed autoencoder-based classification head comprises three key components: (1) an encoder, consisting of convolutional layers followed by fully connected layers, which extracts compact and discriminative feature representations from $F^{\mathrm{LN}}$; (2) a decoder, symmetrically constructed with convolutional and fully connected layers, which reconstructs the input from the latent space to promote feature consistency and preservation; and (3) a classification head, which operates on the latent vector to perform the final classification. The autoencoder is trained with a dual-objective loss that combines reconstruction loss and classification loss. The reconstruction loss ensures that the latent representation retains critical information by minimizing the discrepancy between the original input $F^{\mathrm{LN}}$ and its reconstruction $\hat{F}^{\mathrm{LN}}$:
@@ III-F Stage 4: Retrain Classification Head Based on Autoencoder Network
The classification loss ensures accurate predictions by aligning the predicted probability distribution $\hat{y}_{b}^{i}$ with the ground-truth label $y_{b}^{i}$.
@@
To jointly optimize both the reconstruction and classification objectives, we adopt a task uncertainty weighting strategy [37], which adaptively adjusts the contribution of each loss term based on its inherent uncertainty. The total loss function is expressed as:
@@
where $\sigma_{\text{recon}}$ and $\sigma_{\text{ce}}$ are learnable parameters representing the task-dependent uncertainties for the reconstruction and classification losses, respectively. This formulation allows the model to balance the learning dynamics of both tasks in a principled and data-driven manner.
To control the false alarm rate during inference, we sort the first element in the softmax output for all test clutter samples in descending order, yielding a one-dimensional array $O_{\mathrm{sorted}}$. The detection threshold $\eta$ is then determined based on the desired false alarm rate $P_{fa}^{\text{d}}$ as follows:
@@
where $O_{\mathrm{sorted}}(x)$ denotes the $x$-th largest output in the sorted array, $N_{\text{clutter}}$ is the total number of test clutter samples, and $\lceil\cdot\rceil$ is the ceiling operation to ensure integer indexing.
Figure 6: Overview of the autoencoder model.
@@ IV Experiments
In this section, we first describe the experimental setup, followed by a comprehensive evaluation of RadarLLM’s detection performance under three settings: sufficient training data, constrained training data, and zero-shot scenarios.
@@ IV-A1 Dataset
We employ eight benchmark data sequences (Data1–Data8) from the Intelligent PIxel Processing X-band (IPIX) radar archive, summarized in Table I. All datasets were recorded in 1993 by the IPIX radar on Canada’s east coast and include full polarization measurements (HH, HV, VH, and VV). Each set contains 14 range cells sampled at 1 kHz, with 131 072 complex echoes per cell, yielding an observation window of 0.512 s per sample. Returns from the primary range cell constitute target echoes, whereas returns from the remaining clutter-only cells represent sea clutter. To facilitate performance analysis across varying operating conditions, Data1–Data3 correspond to low signal-to-clutter ratio (SCR) scenes, while Data4–Data8 represent high-SCR scenes.
To ensure sufficient training data, we employ overlapped segmentation following the partition rule in Eq. (1), with parameters set to $M=32$ for target cells and $M=128$ for clutter cells. This process generates 4,079 target samples and more than 9,000 clutter samples per dataset. The samples are divided into three groups: (1) a training set using the first 20% of observation time for both the target and clutter cells, (2) a validation set covering 20% to 35% of the observation time, and (3) a test set containing the remaining samples.
@@ IV-A2 Baselines
  • Bi-LSTM [14]: Bi-LSTM enhances the traditional Long Short-Term Memory (LSTM) by incorporating bidirectional processing. In our implementation, the number of LSTM layers and the number of hidden units per layer are set to 2 and 6, respectively.
@@
  • Transformer [39]: Transformer evolutionizes sequence processing through self-attention mechanisms. In experiments, the Transformer encoder is configured with a model dimensionality of 32, comprising 2 layers and 8 attention heads. The feed-forward network (FFN) within each layer is designed with a hidden size of 128.
@@ IV-A3 Evaluation Metrics
In maritime target detection, achieving a high detection rate at an extremely low false alarm rate is imperative, as misinterpreting sea clutter as targets can have significant consequences, particularly in military operations. We evaluate the detection performance of different detectors at a false alarm rate of 0.005, considering our test set only comprises 6,000 clutter samples. Selecting a lower false alarm rate could result in unstable comparisons due to the limited number of clutter samples.
@@ IV-B Comparison of Detection Performance, Speed, and Network Parameters Under Sufficient Training Samples
Table IV reports the detection rates achieved by RadarLLM and the eight baseline detectors introduced in Section IV-A2. All results are obtained with a fixed false alarm rate of 0.005 on the eight IPIX datasets. Compared to the best competing sequence feature detector, RadarLLM improves the detection rate by 7.60%, 3.85%, 2.30%, and 1.93% for the HH, HV, VH, and VV polarizations, respectively. Figure IV presents the complete ROC curves; it is evident that RadarLLM maintains the highest detection probability across the full range of false alarm rates. We also compare RadarLLM with ADN18, which relies on high-dimensional time–frequency maps and an enhanced CNN backbone. Despite using only compact sequence features, RadarLLM still delivers a 1.14% gain in average detection rate.
Furthermore, we compare RadarLLM with other baseline models in terms of training complexity, total network parameters, and inference cost to evaluate its practicality for real-world deployment. The comparison results are summarized in Table V. Although RadarLLM has the highest total network parameters, the number of trainable parameters is comparable to those of other models, as the majority of its weights are frozen during fine-tuning. Notably, RadarLLM achieves an exceptionally low average inference latency. Specifically, it requires only 5.6561 seconds to process over 9,000 test samples, making it 28.44 times faster than the ADN18 method. This efficiency is primarily attributed to our patching strategy, which substantially reduces the number of feature tokens, and consequently, the computational overhead. Moreover, the intrinsic inference acceleration of the GPT architecture further enhances this performance gain. Together, these advantages make RadarLLM a promising solution for real-time marine target detection.
@@ IV-C Comparison of Detection Performance Under Limited Training Samples With Baselines
Due to the inherent sparsity of radar targets in real-world scenarios, obtaining high-quality labeled samples for radar target detection presents a substantial challenge. Considering this issue, we conduct experiments under constrained training sample conditions to assess the effectiveness of RadarLLM. The samples are divided into three groups: (1) a training set using the first 10% of the observation time for both the target and clutter cells, (2) a validation set covering 10% to 15% of the observation time, and (3) a test set containing the remaining samples.
@@
  • Second, the substantial parameter scale of RadarLLM inherently grants robustness to distributional variations, enabling superior generalization compared to smaller-scale neural networks. Its extensive capacity allows it to capture complex data structures and efficiently handle variability across different radar detection scenarios.
  • Third, the integration of a preference-aware loss function notably improves RadarLLM’s detection performance, especially in challenging scenarios characterized by low signal-to-clutter ratios (SCR).
@@ IV-D Ablation Study
IV-D1 Ablation Study on the Backbone Choice
To validate the effectiveness of using pre-trained LLM, the ablation study is conducted in this section by removing or replacing the pre-trained LLM:
@@ IV-D1 Ablation Study on the Backbone Choice
In Table VII, we show the results of the ablation study. It can be seen that the removal or exchange of pre-trained LLM results in notable performance degradation, indicating the necessity of pretrained LLM for high detection performance.
IV-D2 Ablation Study on the Loss Function
In addition, we compare the fine-tuning effects of preference-aware loss and CE loss on low SCR datasets. First, to facilitate a direct and quantitative comparison between the two loss functions, we aggregate the outputs of all feature tokens by applying a softmax normalization function followed by averaging operation. The resulting probabilities are then used for target detection. In Fig. 9, we evaluate the detection rate on the test samples of different low SCR datasets at different training epochs. It can be seen that preference-aware loss achieves an improvement of around 1%-15% compared to cross-entropy loss (uniformly optimized on all feature patches) in different training epochs and datasets. In Table VIII, we show the detection performance of the reference model. It can be seen that the optimal detection performance of the fine-tuned model outperforms the reference model by approximately 3%-18% across various datasets. Notably, despite this remarkably significant performance gap, leveraging the reference model to score individual feature tokens still yields significant performance gains.
Although the preceding results are promising, a direct comparison between the proposed preference-aware (PA) loss and the conventional end to end cross-entropy (CE) loss remains lacking. To bridge this gap, we evaluate multiple model variants: RadarLLM(PA) vs. RadarLLM(CE), and PatchTST(PA) vs. PatchTST(CE). Models labeled with (PA) adopt the proposed token-level reweighting strategy based on learning value, while those with (CE) rely on standard cross-entropy loss. For further comparison, we include a sample-level reweighting baseline, denoted as WCE [28], where training weights are adjusted at the sample level.
As summarized in Table IX, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions. For instance, RadarLLM(PA) achieves an average gain of 9.9% over RadarLLM(CE), while PatchTST(PA) improves by 7.0% over its CE-based counterpart. Notably, token-level reweighting in both PA variants outperforms WCE, highlighting the advantage of fine-grained reweighting in enhancing detection under challenging scenarios.
@@ IV-D2 Ablation Study on the Loss Function
TABLE VIII: Detection performance of reference model on three low SCR datasets.
@@
TABLE IX: Results of ablation study on the loss function.
@@ IV-E Analysis of the Effect of Hyperparameters on Experimental Results
TABLE X: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarGPT with different number of patch size.)
@@
We first investigate the impact of patch size on RadarLLM’s performance. The results of the ablation study are summarized in Table X. Using a patch size of 48 achieves the highest detection rate, outperforming configurations with smaller (32) and larger (64) patch sizes. These results highlight the importance of selecting an appropriate patch size to optimize the performance of RadarLLM.
We then investigate the impact of the number of GPT2 layers on RadarLLM’s performance. Our ablation study on GPT2 layer configurations is presented in Table XI. Activating four GPT2 layers achieves the highest detection rate, outperforming both shallower configurations and deeper ones. These results highlight two key observations: (1) a model with insufficient layers fails to fully leverage the cross-modal transfer capabilities of the language model, while excess layers suffer from overfitting due to the task-irrelevant network parameters in deeper layers. (2) Although utilizing more GPT2 layers significantly increases the scale of model parameters, the impact on inference latency is minimal. Specifically, the inference speed of GPT2(8) remains approximately 91% of that of GPT2(0), as the inherent inference acceleration of the GPT architecture mitigates the computational overhead introduced by deeper layers.
TABLE XI: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarLLM with different number of GPT2 layers. (RadarLLM(n) represents RadarLLM with n LLM layers.))
@@ IV-F Case Study of the Effectiveness of Feature Token Importance
The SMS feature exhibits a distinctive rise and fall pattern indicative of target presence, primarily characterized by feature tokens 26 and 27. As illustrated in Fig. 9, these tokens initially receive high attention weights during training, underscoring the model’s emphasis on learning these salient patterns. However, as training progresses, the model rapidly assimilates these features, resulting in a marked decline in their weights after approximately 10 epochs. This adaptive behavior reflects the model’s ability to dynamically reallocate attention from well-learned patterns to more complex or informative features, thus mitigating the risk of overfitting to easily recognizable signal features.

### v2

@@ III-D2 Theoretical Insight of the Effectiveness of Feature Token Importance
III-D3 Comparison with Existing Loss Designs
Most existing loss improvements for marine radar target detection address issues such as sample imbalance, hard sample learning, or false alarm control. Various solutions have been proposed, including enhanced focal loss that emphasizes learning from hard or moderately difficult samples [38], and Neyman-Pearson criterion-inspired loss that effectively regulates the false alarm rate [39]. However, these approaches seldom address a crucial issue: mitigating model overfitting in complex real-world scenarios, particularly under low SCR conditions. In contrast, the proposed preference-aware loss incorporates a selective training mechanism that directs the model’s attention to transferable and informative patterns, thus substantially reducing the risk of overfitting. Moreover, while sample reweighting strategies are often employed to alleviate overfitting by defining a weighting function that maps the training loss to a corresponding sample weight, our method adopts a patch-level reweighting strategy motivated by the uneven distribution of informative patterns across multi-domain radar features. This design enables a more fine-grained and flexible learning process. In general, the proposed preference-aware loss marks a major and substantial advance in enhancing model generalization across diverse and low-SCR radar detection scenarios.
III-E False Alarm Controllable Classification Head Retraining
As shown in Fig. 1, the LLM-based model outputs $[B,K,2]$ token-level predictions, which lack a unified detection representation. To achieve consistent decision-making, we retrain a classifier using the learned feature $F^{\mathrm{LN}}$. An autoencoder-based classification head extracts informative features from this high-dimensional input for robust detection. It comprises: (1) an encoder with convolutional and fully connected layers producing compact discriminative features; (2) a symmetric decoder ensuring feature consistency through reconstruction; and (3) a classification head operating on the latent vector. The autoencoder is optimized by a dual-objective loss combining reconstruction and classification terms:
@@ III-E False Alarm Controllable Classification Head Retraining
A task-uncertainty weighting scheme [40] adaptively balances these objectives:
@@
where $\sigma_{\text{recon}}$ and $\sigma_{\text{ce}}$ are learnable task uncertainties that regulate the joint optimization dynamics.
During inference, false alarms are controlled via threshold-setting using sorted softmax outputs of clutter samples. The detection threshold $\eta$ is then determined based on the desired false alarm rate $P_{fa}^{\text{d}}$ as follows:
@@
where $O_{\mathrm{sorted}}(x)$ is the $x$-th largest output and $N_{\text{clutter}}$ is the total number of test clutter samples.
@@ IV Experiments
In this section, we first describe the experimental setup, followed by a comprehensive evaluation of RadarLLM’s detection performance under two settings: sufficient training data and constrained training data scenarios.
@@ IV-A1 Dataset
We employ eight benchmark data sequences (Data1–Data8) from the Intelligent PIxel Processing X-band (IPIX) radar archive, summarized in Table I. All datasets were recorded in 1993 by the IPIX radar on the east coast of Canada and include full polarization measurements (HH, HV, VH, and VV). Each set contains 14 range cells sampled at 1 kHz, with 131 072 complex echoes per cell, yielding an observation window of 0.512 s per sample. Returns from the primary range cell constitute target echoes, whereas returns from the remaining clutter-only cells represent sea clutter. To facilitate performance analysis in varying operating conditions, Data1–Data3 correspond to low signal-to-clutter ratio (SCR) scenes, while Data4–Data8 represent high-SCR scenes.
To ensure sufficient training data, we employ overlapped segmentation following the partition rule in Eq. (1), with parameters set to $M=32$ for target cells and $M=128$ for clutter cells. This process generates 4,079 target samples and more than 9,000 clutter samples per dataset. The samples are divided into three groups: (1) a training set using the first 20% observation time for both the target and clutter cells, (2) a validation set covering 20% to 35% of the observation time, and (3) a test set containing the remaining samples.
@@ IV-A2 Baselines
  • Bi-LSTM [14]: Bi-LSTM enhances traditional long-short-term memory (LSTM) by incorporating bidirectional processing. In our implementation, the number of LSTM layers and the number of hidden units per layer are set to 2 and 6, respectively.
@@
  • Transformer [36]: Transformer reshapes sequence processing through self-attention mechanisms. In experiments, the Transformer encoder is configured with a model dimensionality of 32, comprising 2 layers and 8 attention heads. The feed-forward network (FFN) within each layer is designed with a hidden size of 128.
@@ IV-A3 Evaluation Metrics
In maritime target detection, achieving a high detection rate at an extremely low false alarm rate is imperative, as misinterpreting sea clutter as targets can have significant consequences, particularly in military operations. We evaluated the detection performance of different detectors at a false alarm rate of 0.005, considering that our test set only comprises 6,000 clutter samples. Selecting a lower false alarm rate could result in unstable comparisons due to the limited number of clutter samples.
@@ IV-B Comparison of Detection Performance, Speed, and Network Parameters Under Sufficient Training Samples
Table IV reports the detection rates achieved by RadarLLM and the eight baseline detectors introduced in Section IV-A2. All results are obtained with a fixed false alarm rate of 0.005 on the eight IPIX datasets. Compared to the best competing sequence feature detector, RadarLLM improves the detection rate by 7.60%, 3.85%, 2.30%, and 1.93% for the HH, HV, VH, and VV polarizations, respectively. Figure IV presents the complete ROC curves; it is evident that RadarLLM maintains the highest detection probability across the full range of false alarm rates. We also compare RadarLLM with ADN18, which relies on high-dimensional time–frequency maps and an enhanced CNN backbone. Despite using only compact sequence features, RadarLLM still delivers a 1.14% gain in the average detection rate.
Furthermore, we compare RadarLLM with other baseline models in terms of training complexity, total network parameters, and inference cost to evaluate its practicality for real-world deployment. The comparison results are summarized in Table V. Although RadarLLM has the highest total network parameters, the number of trainable parameters is comparable to those of other models, as the majority of its weights are frozen during fine-tuning. In particular, RadarLLM achieves an exceptionally low average inference latency. Specifically, it requires only 5.6561 seconds to process more than 9,000 test samples, which is 28.44 times faster than the ADN18 method. This efficiency is primarily attributed to our patching strategy, which substantially reduces the number of feature tokens, and consequently, the computational overhead. Moreover, the intrinsic inference acceleration of the GPT architecture further enhances this performance gain. Together, these advantages make RadarLLM a promising solution for real-time marine target detection.
@@ IV-C Comparison of Detection Performance Under Limited Training Samples With Baselines
Due to the inherent sparsity of radar targets in real-world scenarios, obtaining high-quality labeled samples for radar target detection presents a substantial challenge. Considering this issue, we conducted experiments under constrained training sample conditions to assess the effectiveness of RadarLLM. The samples are divided into three groups: (1) a training set using the first 10% of the observation time for both the target and clutter cells, (2) a validation set covering 10% to 15% of the observation time, and (3) a test set containing the remaining samples.
@@
  • Second, the substantial parameter scale of RadarLLM inherently grants robustness to distributional variations, enabling superior generalization compared to smaller-scale neural networks. Its extensive capacity allows it to capture complex data structures and efficiently handle variability in different radar detection scenarios.
  • Third, the integration of a preference-aware loss function significantly improves RadarLLM detection performance, especially in challenging scenarios characterized by low signal-to-clutter ratios (SCR).
@@ IV-D Ablation Study
IV-D1 Ablation Study on Backbone Architectures and Fine-tuning Strategies
To validate the effectiveness of the backbone choice, the ablation study is conducted in this section by removing or replacing the pre-trained LLM:
@@ IV-D1 Ablation Study on Backbone Architectures and Fine-tuning Strategies
In Table VII, we show the results of the ablation study on backbone choice. It can be seen that the removal or exchange of pre-trained LLM results in notable performance degradation, indicating the necessity of pretrained LLM for high detection performance.
@@
Moreover, to verify the effectiveness of the proposed fine-tuning strategy, we conducted an ablation study as summarized in Table VIII, and the corresponding results are presented in Table IX. The results demonstrate that jointly fine-tuning the LoRA modules and Layer Normalization parameters achieves the highest overall performance. In contrast, fine-tuning only the Layer Normalization parameters, the LoRA modules, or performing full fine-tuning results in moderate performance degradation. These findings indicate that the proposed lightweight fine-tuning strategy effectively balances model adaptability and computational efficiency while preserving the universal knowledge embedded in the LLM backbone.
TABLE VIII: Ablation Study on Fine-tuning Strategies settings.
LoRA Layer Norm Other Params
RadarLLM $\checkmark$ $\checkmark$
RadarLLM (F) $\checkmark$ $\checkmark$
RadarLLM (LoRA) $\checkmark$
RadarLLM (LN) $\checkmark$
TABLE IX: Results of ablation study on the finetuning strategy.
$H$ HV VH $V$ AVG
RadarLLM 82.51 84.81 84.61 75.90 81.96
RadarLLM (F) 81.69 83.45 84.41 74.07 80.91
RadarLLM (LoRA) 82.20 84.45 84.56 75.53 81.69
RadarLLM (LN) 82.33 83.73 83.80 75.12 81.25
IV-D2 Ablation Study on the Loss Function
In addition, we compare the fine-tuning effects of preference-aware loss and CE loss on low SCR datasets. First, to facilitate a direct and quantitative comparison between the two loss functions, we aggregated the outputs of all feature tokens by applying a softmax normalization function followed by averaging operation. The resulting probabilities are then used for target detection. In Fig. 7, we evaluate the detection rate on the test samples of different low SCR datasets at different training epochs. It can be seen that preference-aware loss achieves an improvement of around 1%-15% compared to cross-entropy loss (uniformly optimized on all feature patches) in different training epochs and datasets. In Table X, we show the detection performance of the reference model. It can be seen that the optimal detection performance of the fine-tuned model outperforms the reference model by approximately 3%-18% across various datasets. In particular, despite this remarkably significant performance gap, leveraging the reference model to score individual feature tokens still yields significant performance gains.
Although the preceding results are promising, a direct comparison between the proposed preference-aware (PA) loss and the conventional end-to-end cross-entropy (CE) loss remains lacking. To bridge this gap, we evaluated multiple model variants: RadarLLM (PA) vs. RadarLLM (CE), and PatchTST (PA) vs. PatchTST (CE). Models labeled with (PA) adopt the proposed token-level reweighting strategy based on learning value, whereas those labeled with (CE) rely on standard cross-entropy loss. For further comparison, we include a sample-level reweighting baseline, denoted as WCE [27], where training weights are adjusted at the sample level.
As summarized in Table XI, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions. For example, RadarLLM (PA) achieves an average gain of 9.9% over RadarLLM (CE), while PatchTST (PA) improves by 7.0% over its CE-based counterpart. In particular, token-level reweighting in both PA variants outperforms WCE, highlighting the advantage of fine-grained reweighting in enhancing detection under challenging scenarios.
TABLE X: Detection performance of reference model on three low SCR datasets.
@@ IV-D2 Ablation Study on the Loss Function
TABLE XI: Results of ablation study on the loss function.
@@ IV-E Analysis of the Effect of Hyperparameters on Experimental Results
TABLE XII: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarLLM with different number of patch size.)
@@
We first investigate the impact of patch size on RadarLLM’s performance. The results of the ablation study are summarized in Table XII. Using a patch size of 48 achieves the highest detection rate, outperforming configurations with smaller (32) and larger (64) patch sizes. These results highlight the importance of selecting an appropriate patch size to optimize the performance of RadarLLM.
We then investigated the impact of the number of GPT2 layers on RadarLLM’s performance. Our ablation study on GPT2 layer configurations is presented in Table XIII. Activating four GPT2 layers achieves the highest detection rate, outperforming both shallower configurations and deeper ones. These results highlight two key observations: (1) a model with insufficient layers fails to fully leverage the cross-modal transfer capabilities of the language model, while excess layers suffer from overfitting due to the task-irrelevant network parameters in deeper layers. (2) Although utilizing more GPT2 layers significantly increases the scale of model parameters, the impact on inference latency is minimal. Specifically, the inference speed of GPT2 (8) remains approximately 91% that of GPT2 (0), as the inherent inference acceleration of the GPT architecture mitigates the computational overhead introduced by deeper layers.
TABLE XIII: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarLLM with different number of GPT2 layers. (RadarLLM (n) represents RadarLLM with n LLM layers.))
@@ IV-F Case Study of the Effectiveness of Feature Token Importance
The SMS feature exhibits a distinctive rise and fall pattern indicative of target presence, mainly characterized by the feature tokens 26 and 27. As illustrated in Fig. 7, these tokens initially receive high attention weights during training, underscoring the model’s emphasis on learning these salient patterns. However, as training progresses, the model rapidly assimilates these features, resulting in a marked decline in their weights after approximately 10 epochs. This adaptive behavior reflects the model’s ability to dynamically reallocate attention from well-learned patterns to more complex or informative features, thus mitigating the risk of overfitting to easily recognizable signal features.